# Global magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regime

### v1

@@ 3.3.1 Initial conditions
The initial HD fields are set to obey approximate radial and vertical hydrostatic equilibrium in the disc (after neglecting the Lorentz force) via (e.g. Nelson et al., 2013),
$\rho(R,z)$ $=\rho_{0}\left(\frac{R_{i}}{R_{0}}\right)^{-3/2}\exp\left[\frac{GM}{c_{s}^{2}}\left(\frac{1}{r}-\frac{1}{R}\right)\right],$ (10)
@@
[…]
This setup enforces a constant aspect ratio $\varepsilon$, such that the disc is not flared. We adopt $\varepsilon=0.1$ as a compromise between numerical efficiency and the observed geometrical disc thickness (e.g Avenhaus et al., 2018). Note that the above expressions are not at equilibrium in the corona because of the larger temperature in this region and that the configuration has been chosen to provide consistency with future models that are closed at the poles. The remaining velocity components, $u_{r}$ and $u_{\theta}$, are initialised as slightly sub-sonic white noise in the inner disc, to perturb the initial magnetic field configuration.
The ZNF magnetic field configuration is initialised as a purely toroidal field, confined to the active zone via,
$B_{\phi}(R,z)=$ $\sqrt{\dfrac{2\rho_{0}R_{0}^{3/2}c_{s}^{2}}{\beta_{0}R^{3/2}}}\sin\left[4\pi\left(\dfrac{R}{R_{0}}-1.5\right)\right]$ (12)
$\max\left(1-\dfrac{z^{2}}{4H^{2}},\ 0\right),\;\;\;\;\;\;\;\text{if}\;\;1.5R_{0}<R<4R_{0},$
and otherwise it is set to zero. The initial (minimum) midplane plasma-beta parameter is set to $\beta_{0}=20$. This configuration, inspired by previous global ZNF studies (e.g. Flock et al., 2011; Sorathia et al., 2012; Parkin & Bicknell, 2013; Parkin, 2014), ensures that saturated MRI turbulence can efficiently sustain a dynamo over the simulation timescale. We stress that initially there is no poloidal magnetic field, and that the dead zone is free of magnetic fields.
@@ 3.3.2 Grid, resolution and units
The numerical domain is described here for the fiducial simulation ZNF-FID, with alterations for the other four simulations outlined in Table 3.3. In spherical radius, a logarithmic grid with 512 cells spans $r\in[R_{0},10R_{0}]$, such that the radial cell width scales as $\delta{r}\propto{r}$. The azimuthal domain covers $\phi\in[0,2\pi]$ with 512 uniformly spaced cells. In the meridional direction, 128 uniformly spaced cells span $\theta\in[1.27,1.87]$, corresponding to $z=\pm 4H$ about the midplane. This is extended with two geometrically stretched buffer regions of 32 cells each in $\theta\in[0.90,1.27]$ and $\theta\in[1.87,2.25]$, respectively. Therefore, the resulting computational domain is a spherical wedge extending to $z=\pm 8H$ in the meridional direction, as illustrated by the quarter-azimuthal wedge render shown in Fig. 2.
We define two reference timescales: $t_{\text{in}}=2\pi/\Omega_{K}(R_{\text{in}})$ corresponding to the orbital period at the innermost radius, and $t_{\textrm{DZI}}=2\pi/\Omega_{K}(R_{\textrm{DZI}})=8t_{\text{in}}$, corresponding to the orbital period at the dead–active zone interface. Code units are chosen such that $GM=\rho_{0}=T_{0}=R_{0}=1$. Finally, the scale-free nature of the setup is broken by the introduction of the dead–active zone interface, which imposes a fixed physical location for the sharp ionisation transition. Consequently, the length scales in the results are normalised such that ${R_{\textrm{DZI}}}\sim 1$ au. However, it is emphasised that $R_{0}=1$ for the purposes of interpreting the analysis of the fields, which are presented in code units.
@@ 3.3.3 Boundary and internal conditions
The inner radial boundary condition copies $\rho$, $P$, $u_{\theta}$, $B_{\theta}$ and $B_{\phi}$ from the innermost active cell into the ghost cells, whilst $u_{\phi}=R$ enforces solid-body rotation. To prevent inflow from the boundary, $u_{r}$ is copied into the ghost cells if $u_{r}<0$ in the first active cell; otherwise, it is symmetric across the boundary. Azimuthal boundaries are periodic, whilst outflow conditions are applied at the outer radial and meridional boundaries.
These boundary conditions are completed by a buffer zone near the inner radius ($r<1.5$) to mitigate boundary artefacts (Dzyurkevich et al., 2010). Within this region, $\mathbf{u}$ and $\rho$ are relaxed back to their initial values on a timescale of 0.1 local orbits. In addition, an Ohmic damping term, $\text{R}_{m}=0.05\varepsilon^{2}\max(1.5-r,0)$, is applied to suppress magnetic fluctuations (e.g. Fromang & Nelson, 2006). Finally, to avoid restrictive time steps in the low-density corona, a density floor of $\rho_{\text{min}}=10^{-9}$ is imposed, and the Alfvén speed is limited to $v_{\text{A}}\leq 10$, by adjusting the local density rather than the magnetic field.
Figure 2: Snapshot of ZNF-FID at $t_{\textrm{in}}=30$, showing a slice through the disc midplane embedded within a quarter-azimuthal wedge of the full domain. The toroidal magnetic field $B_{\phi}$ highlights the self-consistent turbulent magnetic structures, clearly delineating the midplane location of the dead–active zone interface (black line) at $R_{\text{DZI}}$. Over time, a complex laminar $B_{\phi}$ structure extends throughout the entire dead zone (see Fig. 3 and Section 6). The remaining faces of the wedge show the logarithm of the density $\rho$, revealing the vertically stratified disc embedded into a low-density corona.
@@ 3.4.1 Averages and fields
[…]
where the temporal average is computed over the interval $[t_{1},t_{2}]$ with a temporal resolution of $\Delta t=1/\Omega_{K}(R_{\text{in}})\simeq 0.16t_{\text{in}}$, and the meridional integration is taken over $n$ scale heights either side of the midplane.
@@
3.4.2 Asymmetries
To diagnose the non-axisymmetric components, a Fourier decomposition of $Q$ into its azimuthal (or $m$) modes at the midplane is defined as,
@@ 3.4.2 Asymmetries
which is simply a ratio of the vertical component of the vorticity to the surface density. The deviation of the azimuthal velocity from Keplerian flow, $u_{K}=R\Omega_{K}$, is defined as,
$u_{\phi}^{\prime}=u_{\phi}-u_{K}.$ (19)
@@ 3.4.3 Accretion and stresses
Within the turbulent wind driven accretion paradigm outlined in Section 2.2, stresses – both turbulent and laminar – are essential for mediating angular momentum transport, enabling accretion through the disc (see Lesur, 2021a).
[…]
$\mathcal{R}_{r\phi}$ $=\langle\rho\delta{u_{r}}\delta{u_{\phi}}\rangle$ (20)
$\mathcal{M}_{r\phi}$ $=\frac{-\langle B_{r}B_{\phi}\rangle}{4\pi}$ (21)
where $\delta Q$ denotes the fluctuation of $Q$ relative to its spatio-temporal average: $\delta Q=Q-\langle{Q}\rangle$. Following the seminal work of (Shakura & Sunyaev, 1973) these can be written in dimensionless form by introducing an $\alpha$ parameter:
@@
where the subscripts refer to the Reynolds and Maxwell contributions, respectively. To diagnose the relative contributions of laminar and turbulent angular momentum transport at the dead–active zone interface, the Maxwell stress components are further decomposed into coherent and turbulent components as,
$4\pi\,\mathcal{M}_{r\phi}^{\text{coh}}$ $=-\langle{B_{r}}\rangle\langle{B_{\phi}}\rangle$ (23)
$4\pi\,\mathcal{M}_{r\phi}^{\text{turb}}$ $=\langle{-B_{r}B_{\phi}}\rangle+\langle{B_{r}}\rangle\langle{B_{\phi}}\rangle$
$4\pi\,\mathcal{M}_{\theta\phi}^{\text{coh}}$ $=-\langle{B_{\theta}}\rangle\langle{B_{\phi}}\rangle$
@@
Finally, the plasma-beta parameter $\beta$ is defined as the ratio of the gas pressure to the magnetic pressure,
@@
3.4.4 One-dimensional viscous disc diffusion equation
An alternative and simplified approach to modelling accretion is through the one-dimensional viscous diffusion equation, which neglects any wind-driven contribution and does not assume the source of viscosity is turbulence (Lynden-Bell & Pringle, 1974):
$\frac{\partial\Sigma}{\partial t}=\frac{3}{R}\frac{\partial}{\partial R}\left[\sqrt{R}\frac{\partial}{\partial R}\left(\bar{\nu}\Sigma\sqrt{R}\right)\right]$ (26)
where the disc is assumed to be Keplerian, and $\bar{\nu}$ is the mean kinematic viscosity. Within the thin disc approximation, $\bar{\nu}$ can be related to the $\alpha$-prescription via $\bar{\nu}\sim\alpha{c}_{s}H$. The corresponding average viscous torque $\mathcal{G}$ and mass flux $\mathcal{F}$ are given by,
$\mathcal{G}(R,t)$ $=3\pi\bar{\nu}R^{1/2}\Sigma$ (27)
@@ 3.4.4 One-dimensional viscous disc diffusion equation
In the first part of our results, we examine the evolution of the magnetic field in the MRI active zone and its role in shaping the disc’s accretion structure. We begin in Section 4.1 with an overview of the magnetic field evolution across the entire disc, followed by a focused examination of the MRI active zone in Section 4.2 and its vertical structure, in particular, in Section 4.3. Finally, Section 4.4 investigates the midplane radial accretion structure.
@@
Figure 3: Meridional $(R,z)$ plots of the normalised, azimuthally averaged toroidal field $R\langle{B_{\phi}}\rangle_{\phi}$ for three snapshots of ZNF-FID: $t_{\text{in}}=0$ (left), $t_{\text{in}}=200$ (middle) and $t_{\text{in}}=700$ (right). The dashed black lines denote the disc–corona transition at $z=\pm 4H$, and the dead–active zone interface is at $R=1\,\text{au}$. Note the formation of characteristic MRI turbulence in the inner disc, the presence of strong tightly wound magnetic field structures in the MRI active disc surface layers (Section 4.3) and the complex toroidal magnetic field configuration that eventually permeates the entire dead zone (Section 6.1).
@@ 4.1 Overview of magnetic field evolution in the disc
The dynamics in the simulations are primarily governed by the MRI and the disc–corona temperature transition. We begin by presenting a render of the numerical domain at $t_{\text{in}}=30$ for ZNF-FID in Fig. 2, which illustrates the rapid emergence of a self-consistently MRI-turbulent disc structure in the inner region ($R<1\,\mathrm{au}$), with MRI activity suppressed beyond the dead–active zone interface (solid black line).
To illustrate the overall evolution of magnetic structures in the disc, we present snapshots of the normalised azimuthally averaged toroidal field $R\langle B_{\phi}\rangle_{\phi}$ in the meridional plane $(R,z)$, from the initial condition, through to $t_{\text{in}}=700$, in Fig. 3. Radially, a turbulent MRI active region emerges, with the MRI suppressed in the dead zone. This suppression results from the choice of magnetic Reynolds number (right panel of Fig. 1), which lies below the empirical threshold for stratified ZNF MRI stability, $\text{R}_{m}\lesssim 3000$ (Flock et al., 2012). Over time, a complex toroidal field morphology emerges and gradually extends into the dead zone, a process examined in more detail in Section 6. Meanwhile, in the vertical direction, strong toroidal magnetic field structures build up at the disc–corona temperature transition in the MRI active zone. This configuration exerts significant magnetic torques on the disc, driving localised radial surface layer accretion (see Section 4.3). Finally, large-scale coherent poloidal magnetic field loops are generated in the ideal MHD region, spanning the radial and meridional extent of the MRI active zone; we explore this further in Section 4.2.2.

### v3

@@ 3.3.1 Initial conditions
The initial HD fields are set to obey approximate radial and vertical hydrostatic equilibrium in the disc (after neglecting the Lorentz force) via (e.g. nelson_linear_2013),
$\rho(R,z)$ $=\rho_{0}\left(\frac{R_{\text{i}}}{R_{0}}\right)^{-3/2}\exp\left[\frac{GM}{c_{\text{s}}^{2}}\left(\frac{1}{r}-\frac{1}{R}\right)\right],$ (10)
@@
where the midplane gas density profile has been set in a self-similar fashion such that the initial gas surface density $\Sigma=\int^{\infty}_{-\infty}\rho\,\text{d}z\propto R^{-1/2}$, compatible with some observations (see williams_measuring_2016; miotello_setting_2023).
Note that the above expressions are not at equilibrium in the corona because of the larger temperature in this region and that the configuration has been chosen to provide consistency with future models that are closed at the poles. The remaining velocity components, $u_{r}$ and $u_{\theta}$, are initialised with slightly sub-sonic white noise in the inner disc.
The initial ZNF magnetic field configuration is purely toroidal and confined to the active zone via
$B_{\phi}(R,z)=$ $\sqrt{\dfrac{2\rho_{0}R_{0}^{3/2}c_{\text{s}}^{2}}{\beta_{0}R^{3/2}}}\sin\left[4\pi\left(\dfrac{R}{R_{0}}-1.5\right)\right]$ (12)
$\times\max\left(1-\dfrac{z^{2}}{4H^{2}},\ 0\right),\;\;\;\;\;\;\;\text{if}\;\;1.5R_{0}<R<4R_{0},$
and otherwise it is set to zero. The initial (minimum) midplane plasma-beta parameter is set to $\beta_{0}=20$. This configuration, inspired by previous global ZNF studies (e.g. flock_turbulence_2011; sorathia_global_2012; parkin_global_2013; parkin_global_2014), ensures that saturated MRI turbulence can efficiently sustain a dynamo over the simulation timescale. We stress that initially there is no poloidal magnetic field, and that the dead zone is free of magnetic fields.
@@ 3.3.2 Grid, resolution and units
The numerical domain is described here for the fiducial simulation ZNF-FID, with changes made in the other four simulations outlined in Table 3.3. In spherical radius, a logarithmic grid with 512 cells spans $r\in[R_{0},10R_{0}]$, such that the radial cell width scales as $\delta{r}\propto{r}$. The azimuthal domain covers $\phi\in[0,2\pi]$ with 512 uniformly spaced cells. In the meridional direction, 128 uniformly spaced cells span $\theta\in[1.27,1.87]$, corresponding to $z=\pm 4H$ about the midplane. This is extended by two geometrically stretched buffer regions of 32 cells each in $\theta\in[0.90,1.27]$ and $\theta\in[1.87,2.25]$, respectively. Therefore, the resulting computational domain is a spherical wedge extending to $z=\pm 8H$ in the meridional direction, as illustrated by the quarter-azimuthal wedge rendering shown in Fig. 2.
We define two reference timescales: $t_{\text{in}}=2\pi/\Omega_{\text{K}}(R_{0})$ corresponding to the orbital period at the innermost radius, and $t_{\textrm{DZI}}=2\pi/\Omega_{\text{K}}(R_{\textrm{DZI}})=8t_{\text{in}}$, corresponding to the orbital period at the dead–active zone interface. Code units are chosen such that $GM=\rho_{0}=T_{0}=R_{0}=1$.
Finally, the scale-free nature of the setup is broken by the introduction of the dead–active zone interface, which imposes a fixed physical location for the sharp ionisation transition in the inner disc. Consequently, we normalise the length scales in all the results by ${R_{\textrm{DZI}}}=4R_{0}=1$ au, thereby consistently emphasising that this model applies only to the inner disc. [^3] Whilst numerically convenient, the choice of $\varepsilon=0.1$ leads to an effective midplane temperature roughly 2–3 times higher than expected at $R=R_{\text{DZI}}$. However, this has no bearing on our results, given the relative simplicity of our disc model. However, we stress that $R_{0}=1=0.25\,\text{au}$ for the purposes of interpreting the analysis of the fields, which are presented in code units.
@@ 3.3.3 Boundary and internal conditions
At the inner radial boundary condition, $\rho$, $P$, $u_{\theta}$, $B_{\theta}$ and $B_{\phi}$ are copied from the innermost active cell into the ghost cells, whilst $u_{\phi}=R\Omega_{K}(R_{0})$, thereby imposing solid-body rotation at the inner hole of the spherical domain. To prevent inflow from the boundary, $u_{r}$ is copied into the ghost cells if $u_{r}<0$ in the first active cell; otherwise, it is reflected symmetrically across the boundary. At the outer radial boundary and the meridional boundaries, the outflow conditions prescribed in idefix are applied, enforcing zero-gradients for all variables, except that when $u_{r}\geq 0$ in the first active cell, $u_{r}$ is set to zero. Finally, the azimuthal boundaries are periodic.
These boundary conditions are completed by a buffer zone near the inner radius ($r<1.5R_{0}$) to mitigate boundary artefacts (dzyurkevich_trapping_2010). Within this region, $\mathbf{u}$ and $\rho$ are relaxed to their initial values on a timescale of 0.1 local orbital periods. In addition, an Ohmic-resistive damping term, $\eta_{\text{O}}=0.05\varepsilon^{2}\max(1.5R_{0}-r,0)$, is applied to suppress magnetic fluctuations (e.g. fromang_global_2006). Finally, to avoid restrictive time steps in the low-density corona, a density floor of $\rho_{\text{min}}=10^{-9}\rho_{0}$ is imposed, and the Alfvén speed is limited to $v_{\text{A}}\leq 10\,R_{0}\Omega_{K}(R_{0})$, by adjusting the local density rather than the magnetic field.
Figure 2: Snapshot of ZNF-FID at $t_{\textrm{in}}=30$, showing a slice through the disc midplane embedded within a quarter-azimuthal wedge of the full domain. The toroidal magnetic field, $B_{\phi}$, highlights the self-consistent turbulent magnetic structures, clearly delineating the midplane location of the dead–active zone interface (black line) at $R_{\text{DZI}}=4R_{0}=1$ au. Over time, a complex laminar $B_{\phi}$ structure extends throughout the entire dead zone (see Fig. 3 and Section 6). The remaining faces of the wedge show the density $\rho$, revealing the vertically stratified disc embedded into a low-density corona.
@@ 3.4.1 Averages and fields
$\langle Q\rangle_{\phi}$ $=\frac{1}{\phi_{\textrm{max}}}\int^{\phi_{\textrm{max}}}_{0}Q\;d\phi,$ (13)
$\langle Q\rangle_{\phi,t}$ $=\frac{1}{t_{2}-t_{1}}\int^{t_{2}}_{t_{1}}\langle Q\rangle_{\phi}\;dt,$ (14)
$\langle Q\rangle_{\phi,nH}$ $=\int^{\theta(z=-nH)}_{\theta(z=+nH)}r\sin(\theta)\langle Q\rangle_{\phi}\;d\theta,$ (15)
where the temporal average is computed over the interval $t_{1},t_{2}]$ with a resolution of $\Delta t=0.25t_{\text{in}}$, and the meridional integration is taken over $n$ scale heights either side of the midplane.
@@
3.4.2 Azimuthal asymmetries
To diagnose the non-axisymmetric components, a Fourier decomposition of $Q$ into azimuthal ($m$) modes at the midplane is defined as,
@@ 3.4.2 Azimuthal asymmetries
which is simply a ratio of the vertical component of the vorticity to the surface density. The deviation of the azimuthal velocity from Keplerian flow, $u_{\text{K}}=R\Omega_{\text{K}}$, is defined as $u_{\phi}^{\prime}=u_{\phi}-u_{\text{K}}$.
@@ 3.4.3 Accretion and stresses
Within the turbulence and wind-driven accretion paradigms outlined in Section 2.2, stresses – both turbulent and laminar – are essential for facilitating angular momentum transport, enabling accretion through the disc (see lesur_magnetohydrodynamics_2021).
Throughout the rest of this section, $\langle\cdot\rangle$ denotes a generic average, and its specific meaning is clearly stated when used in the results. The total, averaged spherical radial Reynolds and Maxwell stresses (e.g. mishra_strongly_2020) are defined as,
$\mathcal{R}_{r\phi}$ $=\langle\rho\delta{u_{r}}\delta{u_{\phi}}\rangle,$ (19)
$\mathcal{M}_{r\phi}$ $=\frac{-\langle B_{r}B_{\phi}\rangle}{4\pi},$ (20)
where $\delta Q$ denotes the fluctuation of $Q$ relative to its spatio-temporal average: $\delta Q=Q-\langle{Q}\rangle$. Following the seminal work of shakura_black_1973, these can be written in dimensionless form by introducing an $\alpha$ parameter:
@@
where the subscripts refer to the Reynolds and Maxwell contributions, respectively. To diagnose the relative contributions of laminar and turbulent angular momentum transport at the dead–active zone interface, the Maxwell stress components are further decomposed into coherent and turbulent components via
$4\pi\,\mathcal{M}_{r\phi}^{\text{coh}}$ $=-\langle{B_{r}}\rangle\langle{B_{\phi}}\rangle,$ (22)
$4\pi\,\mathcal{M}_{r\phi}^{\text{turb}}$ $=\langle{-B_{r}B_{\phi}}\rangle+\langle{B_{r}}\rangle\langle{B_{\phi}}\rangle,$
$4\pi\,\mathcal{M}_{\theta\phi}^{\text{coh}}$ $=-\langle{B_{\theta}}\rangle\langle{B_{\phi}}\rangle,$
@@
Finally, the plasma-beta parameter, $\beta$, is defined as the ratio of the gas pressure to the magnetic pressure,
@@
3.4.4 One-dimensional viscous-disc diffusion equation
An alternative and simplified approach to modelling accretion is through the one-dimensional viscous-disc diffusion equation, which neglects any wind-driven contribution and does not assume the source of viscosity is turbulence (lynden-bell_evolution_1974):
$\frac{\partial\Sigma}{\partial t}=\frac{3}{R}\frac{\partial}{\partial R}\left[\sqrt{R}\frac{\partial}{\partial R}\left(\bar{\nu}\Sigma\sqrt{R}\right)\right].$ (25)
Here the disc is assumed to be Keplerian, and $\bar{\nu}$ is the mean kinematic viscosity. Within the thin-disc approximation, $\bar{\nu}$ can be related to the $\alpha$-prescription via $\bar{\nu}\!\sim\!\alpha{c}_{\text{s}}H$. The corresponding average viscous torque, $\mathcal{G}$, and mass flux, $\mathcal{F}$, are given by,
$\mathcal{G}(R,t)$ $=3\pi\bar{\nu}R^{1/2}\Sigma,$ (26)
@@ 3.4.4 One-dimensional viscous-disc diffusion equation
In the first part of our results, we examine the evolution of the magnetic field in the MRI-active zone and its role in shaping the disc’s accretion structure. We begin in Section 4.1 with an overview of the magnetic field evolution across the entire disc, followed by a focused examination of the MRI-active zone in Section 4.2 and its vertical structure, in particular, in Section 4.3. Finally, Section 4.4 investigates the midplane radial accretion structure.
@@
Figure 3: Meridional $(R,z)$ plots of the normalised, azimuthally averaged toroidal field, $R\langle{B_{\phi}}\rangle_{\phi}$, for three snapshots of ZNF-FID: $t_{\text{in}}=0$ (left), $t_{\text{in}}=200$ (middle) and $t_{\text{in}}=700$ (right). The dashed black lines denote the disc–corona transition at $z=\pm 4H$, and the dead–active zone interface is at $R=1\,\text{au}$. Note the formation of characteristic MRI-driven turbulence in the inner disc, the presence of strong tightly wound magnetic field structures in the MRI-active disc surface layers (see Section 4.3), and the complex toroidal magnetic field configuration that eventually permeates the entire dead zone (see Section 6.1).
@@ 4.1 Overview of magnetic field evolution in the disc
The dynamics in the simulations are primarily governed by the MRI and the disc–corona temperature transition. We begin by presenting a rendering of the numerical domain at $t_{\text{in}}=30$ for ZNF-FID in Fig. 2, which illustrates the rapid emergence of a self-consistently MRI-turbulent disc in the inner region ($R<1\,\mathrm{au}$), with MRI activity suppressed beyond the dead–active zone interface.
To illustrate the overall evolution of magnetic structures in the disc, we present snapshots of the normalised, azimuthally averaged toroidal field, $R\langle B_{\phi}\rangle_{\phi}$, in the meridional $(R,z)$ plane, from the initial condition through to $t_{\text{in}}=700$, in Fig. 3. Radially, a turbulent MRI-active region emerges, with the MRI suppressed in the dead zone. This suppression results from the choice of magnetic Reynolds number (right panel of Fig. 1), which lies below the empirical threshold for stratified ZNF MRI stability, $\text{R}_{\text{m}}\lesssim 3000$ (flock_turbulence_2012), effectively acting as a numerical switch for MRI activity. Over time, a complex toroidal field morphology emerges and gradually extends into the dead zone, a process examined in more detail in Section 6. Meanwhile, in the vertical direction, strong toroidal magnetic field structures build up at the disc–corona temperature transition in the MRI-active zone. This configuration exerts significant magnetic torques on the disc, driving localised radial surface-layer accretion (see Section 4.3). Finally, large-scale, coherent poloidal magnetic field loops are generated in the ideal-MHD region. These span the radial and meridional extent of the MRI-active zone and are explored further in Section 4.2.2.